Mrs. T walks east on Guernville road at 1m/s, and the bus flies past her headed east at 40m/s. How long will it take for the bus to be 1 kilometer away from her?
1
Expert's answer
2018-10-03T06:48:09-0400
The speed of the bus relative Mrs. T is
v_rel=V-v=40-1=39 m/s.
The time is
t=d/v_rel =1000/39= 25.6 s.